Dichlorodifluoromethane (CCl2F2, 120.91 g/mol) is a refrigerant and aerosol spray propellant that is commercially known as Freon-12. What is the density of CCl2F2 at 878 torr and 32.0°C? a. 5.60 g/L d. 39.9 g/L b. 53.4 g/L e. 5.40 g/L c. 4.18 g/L
Answer #1

Using, ideal gas equation :

Density, d = PM/RT

M = molar mass = 120.91 g/mol

P = 878 torr = 878/760 atm = 1.155 atm (1atm = 760 torr)

R = gas constant = 0.082 Latm/mol.K

T = 32°C = 32+273.15 = 305.15 K

Substituting values, we get :

Density, d = 1.155×120.91/0.082×305.15

= 5.6 g/L